OpenStudy (anonymous):

Hi i need help on double integrals, using the change of order . i dont know how to get the reversed limits. please help. ∫∫ (2x-3y) dx dy + ∫∫ (2x-3y) dx dy first set of limts for dxdy. the inside integral is from 0 to 2x The outside integral is from 0 to 2 Second set of limts for dxdy. the inside integral is from 0 to 4(3-x) The outside integral is from 2 to 3

OpenStudy (anonymous):

this is for the first integral ? the inside integral is from 0 to 2x The outside integral is from 0 to 2 so it should by dydx if so

OpenStudy (anonymous):

Is it possible for you to show me mate. Because if i do it by dydx i get -16 for the first integral, i checked on wolform alpha and its -16/3

OpenStudy (anonymous):

The question states change the order of intregration too?

OpenStudy (anonymous):

the thing i say is that it is not possible to have an integral of dxdy and the limits the inside integral is from 0 to 2x The outside integral is from 0 to 2 because then you will be stuck with a 'x' in the solution

OpenStudy (anonymous):

how come the answeris -16/3 and i get -16 for the first double integral?

OpenStudy (anonymous):

i have left the limits the same.

OpenStudy (anonymous):

you have a region and you chose the limits ?

OpenStudy (anonymous):

what is given in the question

OpenStudy (anonymous):

i got limts 2 to 0 -6x^2 dx which gives -16?

OpenStudy (anonymous):

∫∫ (2x-3y) dx dy if you solve this by the limits that you gave me you would end up with a 'x' in the answer.. please write the question as it is

OpenStudy (anonymous):

iam trying to solve the first set of integrals first, then combine it with the second set but im getting -16, when it should be -16/3 for the first set of double integrals. i need to get the first park right before i proceed to the second set of integrals.

OpenStudy (anonymous):

you see ∫ (2x-3y) dx from 0 to 2x gives : x^2 -3xy plug limits : 4x^2 -6xy now ∫ 4x^2 -6xydy from 0 to 2 gives 4yx^2 -(xy^2)/ 3 plug limits : 8x^2 - (4/3)x

OpenStudy (anonymous):

but there is something wrong with the limits you gave

OpenStudy (anonymous):

Whats wrong with them?

OpenStudy (anonymous):

please write the question again as it is given to you

OpenStudy (anonymous):

∫∫ (2x-3y) dx dy + ∫∫ (2x-3y) dx dy first set of limts for dxdy. the inside integral is from 0 to 2x The outside integral is from 0 to 2 Second set of limts for dxdy. the inside integral is from 0 to 4(3-x) The outside integral is from 2 to

OpenStudy (anonymous):

the limits are not true .. you cant solve it using those limits!

OpenStudy (anonymous):

only if the first integral is dydx instead of dxdy it is possible

OpenStudy (anonymous):

i will attatch a print screen of question.

OpenStudy (anonymous):

only if the first integral is dydx instead of dxdy it is possible

OpenStudy (anonymous):

im having problem attaching the file, im new to do this.. website

OpenStudy (anonymous):

is it possible that the first integral is dydx instead of dxdy ?

OpenStudy (anonymous):

if so : ∫∫ (2x-3y) dy dx using the limits you provided is : 2xy -1.5y^2 from 0 to 2x is : 4x^2 - 6x^2 = -2x^2 ∫ -2x^2dx from 0 to 2 (-2/3)x^3 = -16/3

OpenStudy (anonymous):

i thnk my lecturer did make a mistake on the question so you dont have to change the order?

OpenStudy (anonymous):

if we want to change the order we should be first looking at the region : 

OpenStudy (anonymous):

is it not x=y/2 to 0 to 4 to 0

OpenStudy (anonymous):

if we want to have dxdy instead of dydx we have : look at what x does ? goes from y/2 to 2 and then y goes from 0 to 4

OpenStudy (anonymous):

i got disconnected

OpenStudy (anonymous):

if we want to have dxdy instead of dydx we have : look at what x does ? goes from y/2 to 2 and then y goes from 0 to 4

OpenStudy (anonymous):

yes i finally got part 1, so you can do it from dxdy after all.

OpenStudy (anonymous):

then the integral ∫∫ (2x-3y) dx dy is : x^2 -3xy as x goes from y/2 to 2 is 4 - 6y - y^2/4 + 1.5y^2 = 4-6y +1.25y^2 ∫4-6y +1.25y^2 dy 4y - 3y^2 + 5y^3/12 from 0 to 4 is -16/3 the same as we got

OpenStudy (anonymous):

yes you can but the first integral is dydx .. (you should understand it because of the limits)

OpenStudy (anonymous):

we can change it into dxdy as we did

OpenStudy (anonymous):

im on part 2 the second double integrals, for the x i got x=12-y/4 to 3 to y 4 to 0. is that right

OpenStudy (anonymous):

the answer? or change or limits ?

OpenStudy (anonymous):

for the limits, i got them by a sketch

OpenStudy (anonymous):

i can get the answer by dydx thats easy but dxdy i dont understand the change of limits lol!